2k+4+1+12k=11(k-10)-10(k-5)
Answer by ReadingBoosters(3246)

Starting with
2k + 4 + 1 + 12k = 11(k-10)-10(k-5)
Combine like terms on the left and distribute on the right
14k + 5 = 11k - 110 - 10k + 50
Combine like terms on the right
14k + 5 = k - 60
Group like terms
14k - k = -60 - 5
13k = 55
k = 55/13 = 4 3/13